Planning week 6 - 13 May 2017
This planning week took place during the Hinode eclipse
There were two ARs on the disk at the start of the week: AR 12654 and 12655. The former was the target of the prior week, but we changed to the recently-emerged 12655 as it was showing some B-class flare activity, although this quickly faded. Since there was no other AR, it remained the default AR for the whole week. Due to the lack of activity of 12655, for 9-13 May I just scheduled Atlas_60 every few orbits.
For OP3 we observed a small equatorial coronal hole that was
near a decayed AR. It may be useful for outflow and abundance
The most interesting observations were:
Tbere was little HOP activity this week, with HOP 81 on Sun
& Mon, HOP 336 on Mon, and HOP 130 on Tue.
Saturday 6 May (OP 1)
Atlas_30 was run at 17:05 UT, and then I ran four more Flare_SNS_v2 at 18:43, 20:18, 21:53 and 23:31 UT. IRIS finished at 23:32 UT. **UPDATE: none of these rasters caught a flare.**
Sunday 7 May (OP1)
IUU_SLOT_148X400 was run at 01:02 and 02:41 to search for bursts. There is no IRIS support for this.
07:35 I supported a SST-IRIS coordination by running IUU_SLOT_148X400
for one orbit. I then switched to Cool_loop_stare at 09:16 UT for the
HOP 81 was supported at 10:57 (north pole), and the Hinode
alignment program supported at 18:11 (north) and 20:43 (east).
Monday 8 May (OP1)
IRIS was coordinating with GREGOR and so I inserted Cool_loop_response at 08:14 UT for one orbit.
HOP 81 was supported at 11:28 at the south pole. HOP 336
(bright points) was then supported with four runs of Large_CH_Map at
18:02, 19:43, 21:22 and 23:00 UT.
Tuesday 9 May (OP1)
Atlas_30 was run at disk center on quiet Sun at 00:38 UT.
Tuesday 9 May (OP2)
This OP began with the HOP 130 full Sun scan, which finished
at 17:30. AR 12655 tracking began after this, so I scheduled five
Atlas_60 rasters with the first at 20:20 UT. This may be matched with an IRIS observation (19:59-21:13).
Wednesday 10 May (OP2)
The AR 12655 atlases continued with ones at 02:49, 09:23,
16:03 and 22:35 UT. The latter pointing has a corresponding IRIS raster from 22:10 to 23:52.
Thursday 11 May (OP2)
We switched to quiet disk center for the remainder of the OP
and I scheduled dhb_atlas_120m_30 (spectral atlas) at 00:08, 01:47,
03:25 and 06:42 UT.
Thursday 11 May (OP 3)
I then began tracking of AR 12655 with a set of four Atlas_60
studies, the first at 18:15 UT. There is a corresponding IRIS raster between 17:50 and 18:41.
Friday 12 May (OP 3)
AR 12655 tracking continued with Atlas_60 at 00:50, 07:18,
13:52 and 23:43 UT. An IRIS raster on the AR is available for 23:43 to 00:34 UT.
The Hinode co-alignment was done at the north (18:46) and east
Saturday 13 May (OP 3)
A final Atlas_60 on AR 12655 was done at 06:28 UT.
Page maintained by Dr Peter R Young.